# DAmodel: Hierarchical Bayesian Modelling of DA White Dwarfs for Spectrophotometric Calibration

### v2

@@ 4.2 Spectroscopic Spline Residuals
| WDFS0122-30∗ | 17.661 | 0.001 | 17.981 | -0.0 | 18.447 | 0.001 | 18.916 | -0.005 | 19.299 | 0.005 | 20.671 | -0.004 |
| WDFS0228-08 | 19.509 | -0.0 | 19.698 | 0.003 | 19.805 | -0.002 | 20.158 | -0.001 | 20.483 | 0.003 | 21.708 | -0.014 |
| WDFS0238-36 | 17.782 | -0.001 | 17.959 | -0.0 | 18.082 | 0.001 | 18.427 | 0.0 | 18.743 | -0.002 | 19.954 | 0.002 |
@@
| WDFS0956-38∗ | 17.691 | -0.002 | 17.844 | 0.002 | 17.837 | 0.013 | 18.168 | -0.0 | 18.481 | -0.0 | 19.666 | -0.01 |
| WDFS1024-00 | 18.248 | 0.003 | 18.497 | -0.0 | 18.89 | -0.002 | 19.303 | 0.003 | 19.651 | -0.001 | 20.955 | -0.004 |
| WDFS1055-36∗ | 17.362 | -0.001 | 17.64 | 0.0 | 17.998 | 0.003 | 18.42 | -0.004 | 18.779 | -0.002 | 20.096 | 0.002 |
| WDFS1110-17∗ | 17.032 | -0.001 | 17.34 | -0.001 | 17.847 | 0.007 | 18.304 | -0.003 | 18.674 | 0.0 | 20.017 | 0.003 |
@@
| WDFS1206+02 | 18.232 | -0.002 | 18.468 | 0.003 | 18.661 | -0.003 | 19.049 | 0.0 | 19.395 | 0.001 | 20.667 | -0.001 |
| WDFS1206-27∗ | 15.729 | -0.001 | 16.027 | 0.001 | 16.465 | -0.0 | 16.91 | 0.002 | 17.278 | -0.001 | 18.619 | -0.002 |
| WDFS1214+45∗ | 16.931 | -0.0 | 17.266 | 0.001 | 17.749 | -0.001 | 18.224 | -0.001 | 18.614 | 0.0 | 20.001 | -0.0 |
| WDFS1302+10∗ | 16.178 | 0.0 | 16.507 | 0.0 | 17.025 | -0.001 | 17.502 | -0.001 | 17.888 | 0.001 | 19.269 | 0.001 |
| WDFS1314-03 | 18.251 | -0.003 | 18.578 | 0.002 | 19.084 | 0.003 | 19.557 | -0.0 | 19.94 | 0.0 | 21.311 | -0.022 |
| WDFS1434-28∗ | 17.829 | -0.0 | 17.962 | 0.001 | 17.96 | -0.003 | 18.271 | 0.003 | 18.569 | -0.001 | 19.729 | -0.004 |
@@
| WDFS1930-52∗ | 16.722 | -0.001 | 17.018 | 0.002 | 17.472 | 0.001 | 17.917 | -0.001 | 18.284 | 0.0 | 19.623 | -0.002 |
| WDFS2101-05 | 18.06 | -0.001 | 18.32 | -0.001 | 18.641 | 0.002 | 19.05 | 0.0 | 19.402 | -0.003 | 20.702 | -0.002 |
| WDFS2317-29∗ | 17.886 | 0.002 | 18.129 | -0.001 | 18.34 | -0.002 | 18.739 | -0.002 | 19.089 | 0.0 | 20.373 | 0.012 |
@@ 5 Future Work
An important science case that motivates the need for our all-sky accurate flux standards is SN Ia cosmology. Most recent efforts to constrain our Universe use compilations of SNe from a variety of surveys (Scolnic et al., 2022; Brout et al., 2022; Rubin et al., 2023). An essential part of the analysis is the cross-calibration of the surveys so they are on a consistent flux scale (Scolnic et al., 2015). Popovic et al. (2025, in prep.) are already using our inferred SEDs in their Dovekie cross-calibration codebase. Their work will combine the present day’s largest SNe Ia samples from the Zwicky Transient Facility (Rigault et al., 2025), Dark Energy Survey (des2024) and more, to further constrain the nature of our Universe.
@@ Acknowledgements
The implementation of DAmodel v1.0 is publically available at: https://github.com/benboyd97/DAmodel. All data used, coordinates and our published v1.0 SEDs can be found at: https://zenodo.org/records/14339960. The theoretical DA WD spectra (Hubeny et al., 2021) used to create the underlying grid can be found at: www.as.arizona.edu/~hubeny/pub/DAgrid24.tar.gz. The SEDs used to tie our work to the CALSPEC system (Bohlin et al., 2020) may be found at: https://archive.stsci.edu/hlsps/reference-atlases/cdbs/current_calspec/. The HST/WFC3 UVIS2 and IR filters (Calamida et al., 2022a) used in this work are available at: https://www.stsci.edu/hst/instrumentation/wfc3/performance/throughputs. Data from other surveys used to calculate residuals is published in Axelrod et al. (2023).
@@ Appendix A Published SEDs
| Theoretical DA WD spectra computed with Tlusty v208 (Hubeny et al., 2021) | www.as.arizona.edu/~hubeny/pub/DAgrid24.tar.gz |
@@ Appendix I Photometric Residuals with different surveys
Using our inferred synthetic SEDs for the WDs, we can integrate over various bandpass filters corresponding to different surveys. For any survey that observed a given standard, we can compute the residual between their measurement and our synthetic prediction. We show our residuals for DES in Figure 9, SDSS in Figure 23, PS1 in Figure 24, Gaia in Figure 25 and DECaLS in Figure 26. The data collection process and residual results are discussed in Section 4.4. Since each survey is calibrated independently, a wavelength-dependent bias is expected. Additional biases are anticipated because our SEDs are tied to the new CALSPEC Bohlin et al. (2020) system rather than Bohlin (2014). In the Figures we show that the residuals are significantly smaller, comparable to those in Axelrod et al. (2023), when we tie our network to the the older CALSPEC system (Bohlin, 2014). In Figure 25 we also find smaller residuals when using the same scaled definition of Vega as Gaia (Riello et al., 2021). Despite getting better agreement with other surveys when using the older CALSPEC system, we maintain our decision to publish the network with respect to the more accurate Bohlin et al. (2020) CALSPEC system and encourage other surveys to adopt the same for their absolute calibration. Quantifying these biases is crucial for cross-calibrating surveys and combining samples. In Section 5, we discuss the ongoing application of our SEDs to unify SN Ia samples across different surveys.

### v3

@@ 4.2 Spectroscopic Spline Residuals
| WDFS0122-30∗ | 17.661 | 0.001 | 17.981 | -3.7e-04 | 18.447 | 0.001 | 18.916 | -0.005 | 19.299 | 0.005 | 20.671 | -0.004 |
| WDFS0228-08 | 19.509 | -3.0e-06 | 19.698 | 0.003 | 19.805 | -0.002 | 20.158 | -0.001 | 20.483 | 0.003 | 21.708 | -0.014 |
| WDFS0238-36 | 17.782 | -0.001 | 17.959 | -2.7e-05 | 18.082 | 0.001 | 18.427 | 4.9e-04 | 18.743 | -0.002 | 19.954 | 0.002 |
@@
| WDFS0956-38∗ | 17.691 | -0.002 | 17.844 | 0.002 | 17.837 | 0.013 | 18.168 | -3.9e-04 | 18.481 | -9.0e-06 | 19.666 | -0.01 |
| WDFS1024-00 | 18.248 | 0.003 | 18.497 | -9.0e-06 | 18.89 | -0.002 | 19.303 | 0.003 | 19.651 | -0.001 | 20.955 | -0.004 |
| WDFS1055-36∗ | 17.362 | -0.001 | 17.64 | 2.0e-04 | 17.998 | 0.003 | 18.42 | -0.004 | 18.779 | -0.002 | 20.096 | 0.002 |
| WDFS1110-17∗ | 17.032 | -0.001 | 17.34 | -0.001 | 17.847 | 0.007 | 18.304 | -0.003 | 18.674 | 2.1e-04 | 20.017 | 0.003 |
@@
| WDFS1206+02 | 18.232 | -0.002 | 18.468 | 0.003 | 18.661 | -0.003 | 19.049 | 4.7e-04 | 19.395 | 0.001 | 20.667 | -0.001 |
| WDFS1206-27∗ | 15.729 | -0.001 | 16.027 | 0.001 | 16.465 | -4.3e-04 | 16.91 | 0.002 | 17.278 | -0.001 | 18.619 | -0.002 |
| WDFS1214+45∗ | 16.931 | -3.3e-04 | 17.266 | 0.001 | 17.749 | -0.001 | 18.224 | -0.001 | 18.614 | 4.2e-04 | 20.001 | -3.0e-05 |
| WDFS1302+10∗ | 16.178 | 2.7e-04 | 16.507 | 2.2e-05 | 17.025 | -0.001 | 17.502 | -0.001 | 17.888 | 0.001 | 19.269 | 0.001 |
| WDFS1314-03 | 18.251 | -0.003 | 18.578 | 0.002 | 19.084 | 0.003 | 19.557 | -3.4e-04 | 19.94 | 9.0e-05 | 21.311 | -0.022 |
| WDFS1434-28∗ | 17.829 | -3.4e-04 | 17.962 | 0.001 | 17.96 | -0.003 | 18.271 | 0.003 | 18.569 | -0.001 | 19.729 | -0.004 |
@@
| WDFS1930-52∗ | 16.722 | -0.001 | 17.018 | 0.002 | 17.472 | 0.001 | 17.917 | -0.001 | 18.284 | 1.6e-04 | 19.623 | -0.002 |
| WDFS2101-05 | 18.06 | -0.001 | 18.32 | -0.001 | 18.641 | 0.002 | 19.05 | 3.3e-04 | 19.402 | -0.003 | 20.702 | -0.002 |
| WDFS2317-29∗ | 17.886 | 0.002 | 18.129 | -0.001 | 18.34 | -0.002 | 18.739 | -0.002 | 19.089 | 2.8e-04 | 20.373 | 0.012 |
@@ 5 Future Work
An important science case that motivates the need for our all-sky accurate flux standards is SN Ia cosmology. Most recent efforts to constrain our Universe use compilations of SNe from a variety of surveys (Scolnic et al., 2022; Brout et al., 2022; Rubin et al., 2023). An essential part of the analysis is the cross-calibration of the surveys so they are on a consistent flux scale (Scolnic et al., 2015). Popovic et al. (2025, in prep.) are already using our inferred SEDs in their Dovekie cross-calibration codebase. Their work will combine the present day’s largest SNe Ia samples from the Zwicky Transient Facility (Rigault et al., 2025), Dark Energy Survey (DES Collaboration et al., 2024) and more, to further constrain the nature of our Universe.
@@ Acknowledgements
The implementation of DAmodel v1.0 is publically available at: https://github.com/benboyd97/DAmodel. All data used, coordinates and our published v1.0 SEDs can be found at: https://zenodo.org/records/14339960. The theoretical DA WD spectra (Hubeny et al., 2021) used to create the underlying grid can be found at: www.as.arizona.edu/˜hubeny/pub/DAgrid24.tar.gz. The SEDs used to tie our work to the CALSPEC system (Bohlin et al., 2020) may be found at: https://archive.stsci.edu/hlsps/reference-atlases/cdbs/current_calspec/. The HST/WFC3 UVIS2 and IR filters (Calamida et al., 2022a) used in this work are available at: https://www.stsci.edu/hst/instrumentation/wfc3/performance/throughputs. Data from other surveys used to calculate residuals is published in Axelrod et al. (2023).
@@ Appendix A Published SEDs
| Tables 1-3 and synthetic photometry for other surveys | https://zenodo.org/records/16877348/files/paper_tables.zip |
@@
| Theoretical DA WD spectra computed with Tlusty v208 (Hubeny et al., 2021) | www.as.arizona.edu/˜hubeny/pub/DAgrid24.tar.gz |
@@ Appendix I Photometric Residuals with different surveys
Using our inferred synthetic SEDs for the WDs, we can integrate over various bandpass filters corresponding to different surveys. For any survey that observed a given standard, we can compute the residual between their measurement and our synthetic prediction. We show our residuals for DES in Figure 9, SDSS in Figure 23, PS1 in Figure 24, Gaia in Figure 25 and DECaLS in Figure 26. The data collection process and residual results are discussed in Section 4.4. Since each survey is calibrated independently, a wavelength-dependent bias is expected. Additional biases are anticipated because our SEDs are tied to the new CALSPEC Bohlin et al. (2020) system rather than Bohlin (2014). In the Figures we show that the residuals are significantly smaller, comparable to those in Axelrod et al. (2023), when we tie our network to the the older CALSPEC system (Bohlin, 2014). In Figure 25 we also find smaller residuals when using the same scaled definition of Vega as Gaia (Riello et al., 2021). Despite getting better agreement with other surveys when using the older CALSPEC system, we maintain our decision to publish the network with respect to the more accurate Bohlin et al. (2020) CALSPEC system and encourage other surveys to adopt the same for their absolute calibration. Quantifying these biases is crucial for cross-calibrating surveys and combining samples. In Section 5, we discuss the ongoing application of our SEDs to unify SN Ia samples across different surveys. Synthetic photometry and residuals for these surveys can be found at: https://zenodo.org/records/16877348/files/paper_tables.zip.